OpenStudy (anonymous):

My grandson is about as many days as my son in weeks, and my grandson is as many months as I am in years. My grandson, my son and I together are 120 years. Can you tell me my age in years ?

OpenStudy (black_ninja123):

I think its 40 because 120 divided by 3 is 40

OpenStudy (anonymous):

no I am 72 years old. Let m be my age in years. If s is my son's age in years, then my son is 52s weeks old. If g is my grandson's age in years, then my grandson is 365g days old. Thus, 365g = 52s. Since my grandson is 12g months old, 12g = m. Since my grandson, my son and I together are 120 years, g + s + m = 120. The above system of 3 equations in 3 unknowns (g, s and m) can be solved as follows. m / 12 + 365 m / (52 x 12) + m = 120 or 52 m + 365 m + 624 m = 624 x 120 or m = 624 x 120 / 1041 = 72. So, I am 72 years old

OpenStudy (anonymous):

As they say, beggars can't be choosers, in fact begger take what they can get. A begger on the street can make one cigarette out of every 6 cigarette butts he finds. After one whole day of searching and checking public ashtrays the begger finds a total of 72 cigarette butts. How many cigarettes can he make and smoke from the butts he found?

OpenStudy (anonymous):

14 If the begger can make a whole cigarette from 6 butts then he can make 12 cigarettes from the 72 he found. Once he smokes those, he then will have another 12 butts, which gives him enough to make another 2 cigarettes. A total of 14.